check wheather the orders pair is a solution of the given equation or inequality
1. y=2/3x-4,(10,4)
2. x=-3(-3,1)
3. 5x+y=10,(1,5)

Answer by stanbon(75887)

check wheather the ordered pair is a solution of the given equation or inequality
1. y=2/3x-4,(10,4)
4 = (2/3)10 -4
8 = (2/3)10
false
-----------------
2. x=-3(-3,1)
-3 = -3
true
------------------
3. 5x+y=10,(1,5)
5*1 + 5 = 10
5 + 5 = 10
true
===============
Cheers,
Stan H.
==========